### v1

# Timelike entanglement entropy with gravitational anomalies

Abstract
We study the timelike entanglement entropy (TEE) in two dimensional conformal field theories (CFT) with gravitational anomalies. We employ analytical continuation to compute the timelike entanglement entropy for a pure timelike interval in such CFTs. We find that, unlike the real part, the imaginary part of the TEE displays an asymmetric dependence on the central charges of the left and right moving modes. We propose that the asymmetric dependence on central charges of the imaginary part of the TEE can be used to probe the presence of gravitational anomalies in chiral CFT. Furthermore, we propose a holographic construction to obtain the timelike entanglement entropy from the bulk dual geometries involving topologically massive gravity in AdS3. The holographic results obtained matches exactly with the dual field theory results.
1 Introduction
The study of quantum entanglement in extended many body systems has recently become a crucial area of research yielding insights across condensed matter physics, quantum gravity and black hole physics. A key tool in this investigation is entanglement entropy which quantifies the entanglement in a given bipartite pure states. For two dimensional conformal field theories (CFTs), the replica technique provides a means to calculate this entanglement entropy [1, 2]. However when left and right moving central charges in a CFT become unequal indicating a gravitational anomaly [3], the CFT exhibits a breakdown of stress-energy conservation at the quantum level. This anomaly arising from the non conservation of the stress-energy tensor implies a coordinate dependent description of the theory. The authors [4] obtained the entanglement entropy in two dimensional conformal field theories with gravitational anomalies using the replica technique and showed that it has an additional contribution coming from the Lorentz anomaly. Alternative methods such as the generalized Rindler approach have also been employed to compute entanglement entropy in such CFTs [5]. Subsequent investigations have focused on analyzing other bipartite mixed state entanglement measures such as reflected entropy, entanglement negativity and balanced partial entanglement entropy (BPE) within these anomalous CFTs [6, 7].
In the context of AdS/CFT, [4] provides a bulk description for the holographic entanglement entropy in the presence of a gravitational anomaly where the bulk dual is described by the topologically massive gravity (TMG) in an asymptotically AdS3 spacetime. The action for the TMG-AdS3 consists of usual Einstein-Hilbert term and a gravitational Chern-Simons term which alters the motion of massive spinning particles propagating in the bulk geometry. Their worldlines take the form of ribbons incorporating a normal frame at each point. Due to its explicit dependence on the connection, the Chern-Simons term is not manifestly invariant under diffeomorphism. This absence of invariance directly results in unequal central charges in the dual field theory. The contribution of the Chern-Simons term to entanglement entropy is determined by the boost needed to propagate the normal frame along the worldline. Consequently, the holographic entanglement entropy is given by the sum of the geodesic length (calculated using the RT formula [8, 9]) and the twist of the normal frame along the geodesic [4]. The holographic entanglement entropy can also be determined using the Rindler method as discussed in [10]. Recently, the correction to the holographic entanglement entropy (also BPE) has also been reproduced through the application of newly introduced surface termed inner RT (IRT) surface [11]. See [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] for more development related to TMG-AdS3/CFT2 duality.
Previous research on entanglement entropy has primarily focused on cases where the subsystem in question is a spacelike separated region. However, recent studies have extended this to include timelike scenarios. In this regard, the authors [22, 23] introduced a new entanglement measure for timelike subsystems termed timelike entanglement entropy. It can be defined as the analytic continuation of entanglement entropy of a spacelike subsystem to a timelike one and exhibits complex values in a CFT2. Holographically as shown in [22, 23], the timelike entanglement entropy (TEE) can be obtained by considering combined spacelike and timelike extremal surfaces (geodesics) in the bulk. Specifically, the spacelike extremal surface contributes the real part of TEE while the timelike extremal surface contributes the imaginary part which is obtained by embedding the Poincaré patch in the global patch. Moreover it has been argued that TEE is a specific example of pseudo entropy [24, 25, 26]. Since there can be many multiple combined spacelike and timelike surfaces homologous to the boundary subsystem which gives infinitely many different complex valued areas, [27] introduced the complex-valued weak extremal surface (CWES) criteria which is a complex valued generalization of the Ryu-Takayanagi extremal surface. The unique surface is then selected by choosing the CWES with the lowest imaginary area. On the other hand, for geometries like Lifshitz spacetime where a global embedding is not known, the holographic computation for TEE must be performed entirely within the Poincaré patch. As demonstrated in [28] by extremization of length functional in Lifshitz metric, the holographic TEE involves two spacelike geodesics connecting the timelike interval endpoints to infinity and a timelike geodesic linking the spacelike geodesic endpoints. The timelike geodesic is smoothly joined with the spacelike geodesics at infinity. This method also reproduces the TEE results in AdS3/CFT2 case [28]. Subsequently this approach was extended to holographic TEE in non-conformal theories and non-relativistic theories [29, 30]. In [31], the authors obtain the holographic TEE by suitably generalizing the Ryu-Takayanagi formula to a complexified bulk geometry (refer to [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] for recent progress). Very recently, a novel generalization of the density matrix to capture correlations between timelike separated subsystems have been introduced [62] and entanglement measures for timelike separated subsystems known as entanglement in time was proposed. They further demonstrated that entanglement in time coincides with the TEE for relativistic quantum field theories. This equivalence also applies to the anomalous CFT which we study here.
In this paper, we examine how gravitational anomalies affect TEE in two dimensional CFTs considering both field theory and holographic descriptions. In this context, we first compute TEE at zero temperature, finite temperature with angular potential and finite angular potential with zero temperature in a CFT2 with a gravitational anomaly. This involves the analytic continuation of the entanglement entropy of a spacelike interval to a timelike one. The resulting TEE is complex and reveals distinct central charge dependencies in its real and imaginary parts. Its imaginary part depends on a single central charge while the real part on both left and right central charges. We then compute TEE holographically, showing that it involves summing on-shell actions of a massive spinning particle moving on the spacelike and timelike extremal worldlines. Given that the holographic TEE involves both spacelike and timelike geodesics, and the Chern-Simons contribution for spacelike geodesics is known from [4], we here derive the corresponding contribution for timelike geodesics. The key difference is that the Chern-Simons contribution now quantifies the rotation of normal frames along the timelike geodesic whereas it measures the twist (or Lorentz boost) of normal vectors for spacelike geodesics. This contribution from the Chern-Simons term exhibits a complex nature similar to the length of a timelike geodesic. Since the intersection of spacelike and timelike geodesics lie deep in the bulk, we prescribe a specific boundary condition for normal vectors which is based on the time direction in the boundary. Further these boundary conditions are used to obtain the Chern-Simons (anomalous) contribution from each type of geodesic to the TEE. We show that interestingly the bulk holographic results agree exactly with the dual field theory results.
The rest of the paper is organized as follows. In section 2 we briefly review about the entanglement entropy in a CFT2 with gravitational anomaly. In section 3 we compute the TEE of a pure timelike interval using analytical continuation in a CFT2 with a gravitational anomaly. Subsequently in section 4 after reviewing the holographic TMG-AdS3/CFT2 duality, we derive the anomaly contribution due to timelike geodesic and propose a holographic construction for the TEE. We then obtain the holographic TEE in a CFT2 dual to Poincaré AdS3, non-extremal rotating BTZ black hole and extremal rotating BTZ black hole. Finally in section 5 we present our summary and discussions.
2 Entanglement entropy in CFT2 with gravitational anomalies
We first briefly review about the two dimensional CFT with a gravitational anomaly [3, 63]. Such theories are described by two copies of the Virasoro algebra with two unequal central charges $c_{L}$ and $c_{R}$, and there is a breakdown of energy-momentum conservation at the quantum level. The anomaly appears as a gravitational anomaly where the stress tensor is symmetric but not conserved. This non-conservation is captured by the anomalous divergence of the stress tensor as [12]
$\nabla_{\mu}T^{\mu\nu}=\frac{c_{L}-c_{R}}{96\pi}g^{\mu\nu}\epsilon^{\alpha\beta}\partial_{\alpha}\partial_{\rho}\Gamma^{\rho}_{\nu\beta}.$ (2.1)
From a bulk perspective, gravitational anomalies arise because the gravitational Chern-Simons term is only invariant under diffeomorphisms up to a boundary term that leads to a non-zero divergence of the dual field theory stress tensor. Alternatively, the anomaly can also arise from a broken Lorentz symmetry which is anomalies under local frame rotations resulting in a conserved stress tensor but not symmetric. This makes the theory frame dependent. One can move between these two perspectives by incorporating a local counterterm into the CFT generating functional [63, 64].
2.1 Zero temperature
In this subsection we review the entanglement entropy computation for a single interval at zero and finite temperature in a CFT2 with a gravitational anomaly as discussed in [4]. Consider a boosted interval described by $A\equiv[z_{1},z_{2}]=[(x_{1},t_{1}),(x_{2},t_{2})]$ and $B=A^{c}$ represent the rest of the system. This configuration is depicted in fig. 1.
Figure 1: Configuration for a single boosted interval on a complex plane.
The entanglement entropy for a generic spacelike interval $A$ can be expressed in terms of the two point twist field correlators as [1, 2]
$S_{A}=\lim_{n\to 1}\frac{1}{1-n}\log\left<\Phi_{n}(z_{1},\bar{z}_{1})\Phi_{-n}(z_{2},\bar{z}_{2})\right>,$ (2.2)
where $\Phi_{n}(z_{1},\bar{z}_{1})$ and $\Phi_{-n}(z_{2},\bar{z}_{2})$ are the twist and the anti twist fields with conformal dimensions $h_{L}=\frac{c_{L}}{24}\left(n-\frac{1}{n}\right)$ and $h_{R}=\frac{c_{R}}{24}\left(n-\frac{1}{n}\right)$. Since $c_{L}$ and $c_{R}$ are unequal, the twist fields have a non-zero spin $s_{n}$ proportional to the difference in anomaly coefficients $(c_{L}-c_{R})$. The scaling dimension and spin of the twist fields are then given by
$\Delta_{n}=\frac{c_{L}+c_{R}}{24}\left(n-\frac{1}{n}\right)\,\,,\,\,s_{n}=\frac{c_{L}-c_{R}}{24}\left(n-\frac{1}{n}\right).$ (2.3)
On using the two point twist correlator
$\Phi_{n}(z_{1},\bar{z}_{1})\Phi_{-n}(z_{2},\bar{z}_{2})\right>=c_{n}z_{12}^{-2h_{L}}\bar{z}_{12}^{-2h_{R}},$ (2.4)
in (2.2), one can obtain the entanglement entropy for a generic spacelike interval $A$ as
$S_{A}=\frac{c_{L}}{6}\log\left(\frac{z_{12}}{\epsilon}\right)+\frac{c_{R}}{6}\log\left(\frac{\bar{z}_{12}}{\epsilon}\right),$ (2.5)
where $\epsilon$ is a UV cut-off. Now analytically continuing to the Lorentzian signature via $z=x-t$, $\bar{z}=x+t$ and using $z=Re^{i\theta}$, the entanglement entropy can be rewritten as follows [4]
$S_{A}=\frac{c_{L}+c_{R}}{6}\log\left(\frac{R}{\epsilon}\right)-\frac{c_{L}-c_{R}}{6}\kappa.$ (2.6)
where $\kappa$ is the boost parameter and related to rotation angle as $\theta=i\kappa$. It is observed that the second term in the above expression arises due to contribution from the gravitational anomaly. The length $R$ and boost $\kappa$ for the boosted interval $A$ is given in terms of $(t,x)$-coordinates as
$R=\sqrt{x_{12}^{2}-t_{12}^{2}}\,,\qquad\kappa=\tanh^{-1}\left(\frac{t_{12}}{x_{12}}\right).$ (2.7)
In the absence of an anomaly ($c_{L}=c_{R}$), the entanglement entropy (2.6) simplifies to the familiar entanglement entropy of a single interval as described in [1, 2].
2.2 Finite temperature and angular potential
First note that the CFT2 at finite temperature with the gravitational anomaly is described by the grand canonical ensemble. The unequal central charges give rise to a conserved spin angular momentum whose conjugate chemical potential is known as the angular potential. Consider a generic spacelike interval $A\equiv[w_{1},w_{2}]$ in the CFT2 at a finite temperature $T=\beta^{-1}$ and with a non zero chemical potential $\Omega$ for the spin angular momentum. The Euclidean partition function for this finite temperature CFT2 with a angular momentum $J$ as the conserved charge is given by
$Z=\text{Tr}\left(e^{-\beta\,H-\beta\Omega_{E}J}\right),$ (2.8)
where $H$ is the Hamiltonian, $\beta$ is the inverse temperature, $J$ is the angular momentum and $\Omega_{E}=-i\Omega$ is the Euclidean angular velocity. This state exhibits non vanishing energy and momentum expectation values as follows
$H=E_{R}+E_{L}-\frac{c_{L}+c_{R}}{24}\,,\quad J=E_{R}-E_{L}+\frac{c_{L}-c_{R}}{24}.$ (2.9)
The left and right moving inverse temperatures $(\beta_{L},\beta_{R})$ associated with the left and right moving modes in terms of $(\beta,\Omega_{E})$ is given by
$\beta_{L}=\beta(1+i\Omega_{E})\leavevmode\nobreak\ ,\quad\beta_{R}=\beta(1-i\Omega_{E})\leavevmode\nobreak\ .$ (2.10)
The CFT2 for this case lives on a twisted cylinder due to the spin angular momentum and is related to those defined on the complex plane through the following conformal transformation
$z=e^{2\pi w/\beta_{L}},\quad\bar{z}=e^{2\pi\bar{w}/\beta_{R}},$ (2.11)
where $(z,\bar{z})$ and $(w,\bar{w}$) are the coordinates on the complex plane and the twisted cylinder respectively. The entanglement entropy for an interval $A$ may now be obtained by computing the two point twist correlator on the twisted cylinder using the above conformal transformation as
$S_{A}=\frac{c_{L}}{6}\log\left[\frac{\beta_{L}}{\pi\epsilon}\sinh\frac{\pi w_{12}}{\beta_{L}}\right]+\frac{c_{R}}{6}\log\left[\frac{\beta_{R}}{\pi\epsilon}\sinh\frac{\pi\bar{w}_{12}}{\beta_{R}}\right],$ (2.12)
where $w_{12}=x_{12}+i\tau_{12}$ and $\bar{w}_{12}=x_{12}-i\tau_{12}$ are the interval lengths of $A$ on the twisted cylinder. The above expression factorizes into the entanglement entropy for left and right moving temperatures. To see a contribution from the anomaly, on rearranging (2.12) and analytically continuing to Lorentzian signature via $\tau\to it$, the entanglement entropy for an interval $A$ at finite temperature with a conserved angular momentum is given by
$S_{A}=\frac{c_{L}+c_{R}}{12}\log\left[\frac{\beta_{L}\beta_{R}}{\pi^{2}\epsilon^{2}}\sinh\frac{\pi(x_{12}-t_{12})}{\beta_{L}}\sinh\frac{\pi(x_{12}+t_{12})}{\beta_{R}}\right]-\frac{(c_{L}-c_{R})}{12}\log\left[\frac{\beta_{R}\sinh\frac{\pi(x_{12}+t_{12})}{\beta_{R}}}{\beta_{L}\sinh\frac{\pi(x_{12}-t_{12})}{\beta_{L}}}\right].$ (2.13)
Here the second term represents the contribution from gravitational anomaly. The above expression reduces to the entanglement entropy of an interval in a CFT2 at finite temperature with an angular potential [65] in the absence of an anomaly.
2.3 Finite angular potential and zero temperature
Consider a generic spacelike interval $A\equiv[w_{1},w_{2}]$ in a CFT2 at a finite angular potential and zero temperature. For this case one of the inverse temperatures i.e $\beta_{L}$ goes to infinity and the other one $\beta_{R}$ remains finite. The conformal transformation relating correlation functions on the complex plane to those on the twisted cylinder is given by [^1] Since CFT2’s with a conserved angular momentum at zero temperature are dual to the extremal rotating BTZ black holes, the corresponding conformal transformation can be obtained by taking the limit $r\to\infty$ in (4.53) [66].
$z=w,\quad\quad\bar{z}=\frac{\ell}{2r_{0}}e^{2\,r_{0}\bar{w}/\ell}.$ (2.14)
By utilizing the above conformal transformation, the entanglement entropy for an interval $A$ can be obtained by evaluating the two-point twist correlator on the twisted cylinder as [66]
$S_{A}=\frac{c_{L}}{6}\log\frac{w_{12}}{\epsilon}+\frac{c_{R}}{6}\log\left(\frac{\ell}{r_{0}\epsilon}\sinh\frac{r_{0}\bar{w}_{12}}{\ell}\right),$ (2.15)
where $w_{12}=x_{12}+i\tau_{12}$ and $\bar{w}_{12}=x_{12}-i\tau_{12}$. As described in [66], the first part of the above equation is the entanglement entropy of an interval $A$ at zero temperature and the second part is the entanglement entropy at an effective temperature $r_{0}/\pi\ell$. This structure arises because the left moving modes of the CFT are in their ground state while the right moving modes are at an effective temperature known as the Frolov-Thorne temperature $T_{FT}=r_{0}/\pi\ell$ [67]. To see the anomaly contribution explicitly, upon rearranging the above expression and analytically continuing to the Lorentzian signature leads to the entanglement entropy of an interval at zero temperature with finite angular momentum as follows
$S_{A}=\frac{c_{L}+c_{R}}{12}\log\left[\frac{\ell(x_{12}-t_{12})}{r_{0}\epsilon^{2}}\sinh\frac{r_{0}(x_{12}+t_{12})}{\ell}\right]-\frac{(c_{L}-c_{R})}{12}\log\left[\frac{\ell}{r_{0}(x_{12}-t_{12})}\sinh\frac{r_{0}(x_{12}+t_{12})}{\ell}\right].$ (2.16)
The second term in the above expression represents contribution due to the gravitational anomaly.
3 Timelike entanglement entropy in CFT2 with gravitational anomaly
We now proceed to compute the timelike entanglement entropy (TEE) of an interval in a CFT2 with unequal central charges or gravitational anomaly involving different scenarios i.e at zero temperature, finite temperature with angular potential and finite angular potential with zero temperature. This involves the analytic continuation of the entanglement entropy for a spacelike subsystem to a timelike subsystem.
3.1 Zero temperature
The timelike entanglement entropy for a timelike interval $A$ in CFT2 in the presence of gravitational anomaly is obtained by analytically continuing (2.6) to the timelike interval case where $t_{12}^{2}-x_{12}^{2}>0$. On doing so, (2.6) becomes
$S_{A}^{T}=\frac{c_{L}+c_{R}}{6}\log\frac{\sqrt{t_{12}^{2}-x_{12}^{2}}}{\epsilon}+\frac{c_{L}+c_{R}}{12}\log(-1)-\frac{c_{L}-c_{R}}{6}\tanh^{-1}\left(\frac{t_{12}}{x_{12}}\right).$ (3.1)
Here the function $\tanh^{-1}z$ is defined in terms of the complex logarithm as
$\tanh^{-1}z=\frac{1}{2}\log\frac{1+z}{1-z}.$ (3.2)
Now for a pure timelike interval, we set $x_{12}=0$ and $t_{12}=T$ in (3.1) to obtain the timelike entanglement entropy as [^2] Since for a pure timelike interval $\frac{t_{12}}{x_{12}}\to\infty$ as $x_{12}\to 0^{+}$ and $t_{12}>0$, one obtain from (3.2) $\tanh^{-1}\infty=\frac{1}{2}\log(-1).$ (3.3)
$S_{A}^{T}=\frac{c_{L}+c_{R}}{6}\log\frac{T}{\epsilon}+\frac{c_{L}+c_{R}}{12}\log(-1)-\frac{c_{L}-c_{R}}{12}\log(-1).$ (3.4)
We will consider the principal value of log from now on, then the timelike entanglement entropy in the presence of anomaly for a pure timelike interval is given by
$S_{A}^{T}$ $=\frac{c_{L}+c_{R}}{6}\log\frac{T}{\epsilon}+\frac{c_{L}+c_{R}}{12}i\pi-\frac{c_{L}-c_{R}}{12}i\pi$ (3.5a)
$=\frac{c_{L}+c_{R}}{6}\log\frac{T}{\epsilon}+\frac{c_{R}}{6}i\pi.$ (3.5b)
This result (3.5b) generalizes that of [22, 23] for the timelike entanglement entropy in non-anomalous CFT. In the presence of gravitational anomaly, the TEE acquires an additional contribution proportional to the anomaly coefficient $c_{L}-c_{R}$ (third term in (3.5a)) and a partial cancellation of the usual imaginary part of the TEE with the contribution from the gravitational anomaly occurs. This results in an asymmetric dependence of the TEE on the central charges, with the real part depends on both central charges and the imaginary part depends only on one of the two central charges. In particular, the TEE of the left moving mode is purely real. This later property is an useful feature that one may exploit to detect the presence of gravitational anomaly in a chiral CFT.
3.2 Finite temperature and angular potential
We now compute the timelike entanglement entropy in a two dimensional CFT at finite temperature and non-zero chemical potential. The timelike entanglement entropy is obtained by analytically continuing the spacelike interval $A$ in (2.13) to the timelike interval by making $t^{2}_{12}>x^{2}_{12}$ as
$S_{A}^{T}$ $=\frac{c_{L}+c_{R}}{12}\log\left[\frac{\beta_{L}\beta_{R}}{\pi^{2}\epsilon^{2}}\sinh\frac{\pi(t_{12}-x_{12})}{\beta_{L}}\sinh\frac{\pi(t_{12}+x_{12})}{\beta_{R}}\right]-\frac{(c_{L}-c_{R})}{12}\log\left[\frac{\beta_{R}\sinh\frac{\pi(t_{12}+x_{12})}{\beta_{R}}}{\beta_{L}\sinh\frac{\pi(t_{12}-x_{12})}{\beta_{L}}}\right]$ (3.6)
$+\frac{c_{L}+c_{R}}{12}i\pi-\frac{(c_{L}-c_{R})}{12}i\pi.$
Then the timelike entanglement entropy for a pure timelike interval at finite temperature and angular potential can be obtained by taking $x_{12}=0$ and $t_{12}=T$ in the above expression as follows
$S_{A}^{T}=\frac{c_{L}+c_{R}}{12}\log\left[\frac{\beta_{L}\beta_{R}}{\pi^{2}\epsilon^{2}}\sinh\frac{\pi T}{\beta_{L}}\sinh\frac{\pi T}{\beta_{R}}\right]-\frac{(c_{L}-c_{R})}{12}\log\left[\frac{\beta_{R}\sinh\frac{\pi T}{\beta_{R}}}{\beta_{L}\sinh\frac{\pi T}{\beta_{L}}}\right]+\frac{c_{R}}{6}i\pi.$ (3.7)
We find that the above TEE is complex where only one central charge contributes to the imaginary part and both central charges are present in the real part, similar to the zero temperature case. However the anomaly contribution to the real part (proportional to $c_{L}-c_{R}$) is present in contrast to earlier case. It matches with the CFT2 result obtained in [23] when the anomaly is absent i.e. $c_{L}=c_{R}$. We note that in the zero temperature limit $(\beta_{L},\beta_{R})\gg T$, the above expression reduces to (3.5). We also note that one can consider the Rindler approach [43] in the presence of gravitational anomaly and obtain the same results (3.7) and (3.5).
3.3 Zero temperature and finite angular potential
Next, we obtain the timelike entanglement entropy of a pure timelike interval in a CFT2 at zero temperature characterized by finite angular momentum. This can be computed by the analytical continuation of a spacelike interval $A$ to the timelike one followed by taking $x_{12}=0$ and $t_{12}=T$ in (2.16) as follows
$S_{A}^{T}=\frac{c_{L}+c_{R}}{12}\log\left[\frac{\ell\,T}{r_{0}\epsilon^{2}}\sinh\frac{r_{0}T}{\ell}\right]-\frac{(c_{L}-c_{R})}{12}\log\left[\frac{\ell}{r_{0}T}\sinh\frac{r_{0}T}{\ell}\right]+\frac{c_{R}}{6}i\pi.$ (3.8)
We observe that, like all the previous cases, the TEE for this case is complex with only one central charge appearing in the imaginary part and both central charges appearing in the real part.
4 Holographic Timelike entanglement entropy in AdS3/CFT2 with gravitational anomaly
In this section, we compute the timelike entanglement entropy holographically in the context of the AdS/CFT correspondence for dual conformal field theories with a gravitational anomaly. In this case the bulk dual geometry is described by topologically massive gravity (TMG) in a bulk AdS3 spacetime [12, 4].
4.1 Review of earlier works
We first briefly review about the holographic description in the context of Topologically Massive Gravity (TMG) in AdS3. The bulk action for TMG in AdS3 consists of the standard Einstein-Hilbert action and an additional gravitational Chern-Simons (CS) term which is given by [68, 69, 4]
$S_{\rm TMG}={1\over 16\pi G_{N}^{(3)}}\int d^{3}x\sqrt{-g}\left(R+{2\over\ell^{2}}\right)+\frac{1}{32\pi G_{N}^{(3)}\mu}\int d^{3}x\sqrt{-g}\epsilon^{\lambda\mu\nu}\Gamma^{\rho}_{\lambda\sigma}\left(\partial_{\mu}\Gamma^{\sigma}_{\rho\nu}+{2\over 3}\Gamma^{\sigma}_{\mu\tau}\Gamma^{\tau}_{\nu\rho}\right),$ (4.1)
where $\mu$ is a parameter of mass dimension governing the coupling of the CS term and $\ell$ represents the AdS3 radius. The equations of motion derived from this action take the following form
$R_{\mu\nu}-{1\over 2}g_{\mu\nu}R-{1\over\ell^{2}}g_{\mu\nu}=-{1\over\mu}C_{\mu\nu},$ (4.2)
where $C_{\mu\nu}$ denotes the Cotton tensor. When the Cotton tensor vanishes, the theory admits solutions that retain the form of Einstein metrics ensuring that these solutions remain locally AdS3. In this work, we confine our discussion to such locally AdS3 solutions. The asymptotic symmetry analysis of TMG in AdS3 reveals that the algebra of asymptotic Killing vector modes corresponds to two copies of the Virasoro algebra with left and right moving central charges. This results in a Brown-Henneaux [70] type relation for TMG-AdS3 given by [71, 72]
$c_{L}={3\ell\over 2G_{N}^{(3)}}\left(1-{1\over\mu\ell}\right)\leavevmode\nobreak\ ,\quad c_{R}={3\ell\over 2G_{N}^{(3)}}\left(1+{1\over\mu\ell}\right).$ (4.3)
This relation suggests that the dual conformal field theory exhibits a gravitational anomaly.
Holographic entanglement entropy in TMG-AdS3
For AdS/CFT correspondence, the authors [4] proposed a bulk description for the holographic entanglement entropy in the presence of gravitational anomalies by extending the Lewkowycz-Maldacena prescription [73]. They showed that the holographic entanglement entropy in such cases can be obtained by extremizing a worldline action for a spinning particle propagating in the AdS3 bulk. Consequently, the holographic entanglement entropy is given by the on-shell action of a massive spinning particle in the topologically massive gravity as
[…]
where $C$ represents the worldline of the particle. Here $n_{1}$ and $n_{2}$ are the normal vectors to the curve (up to a choice of handedness) such that $n_{1}\cdot n_{2}=0$. The subscript $E$ indicates that the above expression is evaluated in the Euclidean signature. For locally AdS3 spacetimes where the worldline that minimizes the action is a geodesic, the first part in the above equation corresponds to the Ryu-Takayanagi term originating from the Einstein-Hilbert action and the second part arises from the Chern-Simons term.
When the path (curve) $C$ is spacelike, one can choose $n_{1}=\partial_{t}$ and $n_{2}=\partial_{x}$ where they are both spacelike in the Euclidean signature. On analytically continuing to the complex coordinates $(z,\bar{z})$ via $z=x-t,\bar{z}=x+t$, the normal vectors $(n,\tilde{n})$ in Lorentzian signature is given by
$n:=in_{1}=\partial_{t}\qquad\tilde{n}:=n_{2}=\partial_{x}.$ (4.5)
In Lorentzian signature, one of the normal vector becomes timelike $(n)$ and the other remains spacelike $(\tilde{n})$. The entanglement entropy (4.4) in terms of these normal vectors can be written as
[…]
The above expression gives the on-shell action for the case when the path $C$ is spacelike. Here $s$ parametrizes the length along the path (worldline) $C$ of the particle, $n$ and $\tilde{n}$ are unit timelike and spacelike vectors respectively, both normal to the trajectory of the particle $X^{\mu}$. The second term in (4.6) corresponds to the Chern-Simons contributions and denoted by $S_{\rm anom}^{\rm spacelike}$ can be written in terms of a single normal vector $n$ using $\tilde{n}^{\mu}=\epsilon^{\mu\nu\rho}v_{\nu}n_{\rho}$ as
[…]
where $v^{\mu}=\dot{X}^{\mu}$ is the tangent vector to the path $C$. The above integral is insensitive to the variation of $n$ along the path $C$, and depends only on the initial and the final value of the normal vector $n$ ($\tilde{n}$). The boundary values of $n$ are denoted by
$n(s_{i})=n_{i},\quad n(s_{f})=n_{f},$ (4.8)
where $n_{i}$ and $n_{f}$ are normal vectors at the end point of the path (or geodesic). As described in [4], the integral (4.7) measures the twist of $n_{f}$ relative to $n_{i}$ along $C$. To obtain a compact expression for this integral, one can introduce a vector $n(s)$ that obeys the boundary condition (4.8) and a normal frame along the path defined by the pair of vectors $(q^{\mu},\tilde{q}^{\mu})$ which are both parallel transported i.e. $\nabla q=\nabla\tilde{q}=0$. The normal vector $n(s)$ can then be written in terms of $q$ and $\tilde{q}$ as
$n(s)=\cosh(\eta(s))q(s)+\sinh(\eta(s))\tilde{q}(s),$ (4.9)
where $\eta$ is the rapidity of the Lorentz boost required to boost the orthonormal frame $(v,q,\tilde{q})$ to $(v,n,\tilde{n})$. The integral (4.7) then simplifies to a total derivative given by
[…]
The anomalous contribution from the spacelike curve can be written in terms of $q$ and $\tilde{q}$ by utilizing (4.9) and (4.10) as [4]
$S_{\rm anom}^{\rm spacelike}=\frac{1}{4G_{N}^{(3)}\mu}\log\left(\frac{q(s_{f})\cdot n_{f}-\tilde{q}(s_{f})\cdot n_{f}}{q(s_{i})\cdot n_{i}-\tilde{q}(s_{i})\cdot n_{i}}\right).$ (4.11)
Here $s$ parametrizes the spacelike geodesic and the vectors $(q(s),\tilde{q}(s))$ describes a reference parallel transported normal frame satisfying the following condition
$v^{2}=\tilde{q}^{2}=1$ $,\quad q^{2}=-1,\quad q\cdot\tilde{q}=v\cdot q=v\cdot\tilde{q}=0$ (4.12)
$\nabla v=0\ ,\quad\nabla q=0\ ,\quad\nabla\tilde{q}=0.$
The holographic entanglement entropy is determined only by spacelike geodesic in the bulk and hence, it is obtained by extremizing the on-shell action (4.6) as follows [4]
$S_{\rm HEE}=\mathrm{min}_{C}\,\,\mathrm{ext}\,S_{\rm on-shell}^{\rm spacelike}=\mathrm{min}_{C}\,\,\mathrm{ext}\,\,(S_{\rm EH}^{\rm spacelike}+S_{\rm anom}^{\rm spacelike}),$ (4.13)
which is the sum of length of the geodesic coming from the Einstein-Hilbert term and the twist from the Chern-Simons contribution.
4.2 Anomalous contribution from timelike geodesic
Having described the anomalous contribution from the spacelike geodesic to the holographic entanglement entropy, we now proceed to derive the anomalous (or CS) contribution from timelike geodesic to the on-shell action (4.4). For a timelike path (or geodesic) denoted as $C_{\tau}$, the tangent vector is timelike and the other two normal vectors are spacelike which is different from the spacelike geodesic case. We now choose the normal vectors $(n,\tilde{n})$ in (4.4) as
$n:=n_{1}=\partial_{u}\qquad\tilde{n}:=n_{2}=\partial_{x}.$ (4.14)
These vectors remain spacelike under analytic continuation to Lorentzian signature. We now find the following expression for the on-shell action for a timelike curve in terms of these vectors in the Lorentzian signature using (4.4) as
[…]
where $\tau$ parameterizes the timelike curve $C_{\tau}$. We see that (4.15) becomes imaginary as length of timelike curve is imaginary and also anomalous term becomes imaginary. Since the first term in (4.15) can be obtained by computing the geodesic length, we now focus on the second term denoted by $S^{\rm timelike}_{\rm anom}$. It can be written in terms of a single normal vector $n$ using $\tilde{n}^{\mu}=\epsilon^{\mu\nu\rho}v_{\nu}n_{\rho}$ as
[…]
Similar to the earlier section, the integration only depend on the boundary terms i.e. it depends only on the the initial and the final values of the normal vector $n$ (or $\tilde{n}$). We set the boundary values of the normal vector $n(\tau)$ as
$n(\tau_{i})=n_{i},\quad n(\tau_{f})=n_{f},$ (4.17)
where $n_{i}$ and $n_{f}$ are normal vectors located at the end point of the curve. As both the normal vectors ($n,\tilde{n}$) are spacelike for this case, they rotate along the curve as we parallel transport them from initial to the final point. In other words when we parallel transport $n_{i}$ to the final point, the resulting vector is related to $n_{f}$ via $SO(2)$ transformation. [^3] They are related via $SO(1,1)$ transformation for the case of spacelike curves where one normal vector is timelike and other is spacelike [4]. Now proceeding similar to spacelike curve case, consider a vector $n(\tau)$ that satisfies the boundary condition (4.17) and a parallel transported normal frame which consists of two vectors $(q^{\mu},\tilde{q}^{\mu})$ satisfying $\nabla q=\nabla\tilde{q}=0$. We can expand $n(\tau)$ in terms of $q$ and $\tilde{q}$ as
$n(\tau)=\cos(\theta(\tau))q(\tau)-\sin(\theta(\tau))\tilde{q}(\tau).$ (4.18)
Now using (4.18) in (4.16) leads to
[…]
We observe that the anomaly contribution measures the rotation of frames along the timelike curve which is different from spacelike case where it measures the twist or Lorentz boost along the curve. On using (4.18) in (4.19), the anomaly contribution can be obtained in terms of $q$ and $\tilde{q}$ as
$S_{\rm anom}^{\rm timelike}=\frac{1}{4G_{N}^{(3)}\mu}\log\left[\frac{q(\tau_{f})\cdot n_{f}-i\,\tilde{q}(\tau_{f})\cdot n_{f}}{q(\tau_{i})\cdot n_{i}-i\,\tilde{q}(\tau_{i})\cdot n_{i}}\right].$ (4.20)
Although the angle of rotation of normal frames i.e $\int_{C_{\tau}}\tilde{n}\cdot\nabla n$ along the timelike geodesic is real, the overall anomaly contribution is complex for this case. The parallel transported vectors used above satisfy the following condition
$v^{2}=-1$ $,\quad q^{2}=\tilde{q}=1,\quad q\cdot\tilde{q}=v\cdot q=v\cdot\tilde{q}=0$ (4.21)
$\nabla v=0\ ,\quad\nabla q=0\ ,\quad\nabla\tilde{q}=0.$
As holographic TEE involves both spacelike and timelike geodesics, it is determined by the total of the on-shell actions described by (4.6) and (4.15) as follows
$S_{\rm TEE}=\mathrm{ext}_{C}\leavevmode\nobreak\ \left(S_{\rm on-shell}^{\rm spacelike}+S_{\rm on-shell}^{\rm timelike}\right).$ (4.22)
This expression incorporates the geodesic length (RT formula) along with the anomalous contributions arising from Chern-Simons term considering both spacelike and timelike geodesic contributions. We now utilize this to determine the holographic TEE in the following section.
4.3 Holographic TEE at zero temperature
Let us consider a pure timelike interval $A$ of length $T$ described by $A\equiv[-\frac{T}{2},\frac{T}{2}]$ in a dual CFT2 with a gravitational anomaly at zero temperature. The corresponding holographic dual is given by the AdS3 spacetime in Poincaré coordinates as
$ds^{2}={\ell^{2}\over u^{2}}(-dt^{2}+dx^{2}+du^{2}),$ (4.23)
where $\ell$ is the AdS radius. As discussed in [28] for the case of AdS3/CFT2, the geodesics required for computing the holographic TEE consists of two spacelike geodesics connecting the endpoints of $A$ and infinities plus a timelike geodesic which connects the endpoints of two spacelike geodesics. The equation for these geodesics in the Poincaré metric (4.23) is given by
$t^{2}-u^{2}$ $=\frac{T^{2}}{4}$ (4.24a)
$u^{2}-t^{2}$ $=R^{2},$ (4.24b)
where $T$ and $R$ are constants. The curve (4.24a) is spacelike and describes a geodesics that starts from the end point $t=\pm T/2$ and ends at infinity. These spacelike geodesics are smoothly joined at infinity by a timelike geodesic (4.24b) for any $R>0$. It is depicted in fig. 2.
Figure 2: RT surfaces (geodesics) for a pure timelike interval in Poincaré coordinates are labelled by red and green curves. The normal vectors $n_{i}$ and $n_{f}$ are timelike for spacelike geodesics.
In the presence of an anomaly, the holographic TEE acquires an additional contribution due to the Chern-Simons term present in the bulk gravitational action. Since the holographic TEE consists of both spacelike and timelike geodesics, we must compute the contribution from each geodesic separately. This involves computing the geodesic length and the twist (or rotation) of the normal vectors along each type of geodesic as described in the following subsections.
Spacelike geodesic contribution
The equation for the spacelike geodesics as denoted by red color in fig. 2 is given by (4.24a). The total length of these geodesics can be obtained as
$L^{\rm spacelike}_{A}=2T\ell\int_{\epsilon}^{\infty}\frac{du}{2u}\frac{1}{\sqrt{u^{2}+\frac{T^{2}}{4}}}=2\ell\log{\frac{T}{\epsilon}}.$ (4.25)
Now for anomalous contribution, the tangent vector and parallel transported normal frame vectors along this spacelike geodesic is given by (in $(t,x,u)$ convention)
$v^{\mu}=-\frac{2u}{T\ell}(u,0,t),\quad q^{\mu}=-\frac{2u}{T\ell}(t,0,u),\quad\tilde{q}^{\mu}=(0,\frac{u}{\ell},0).$ (4.26)
These vectors satisfy the condition given in (4.12). While the normal vector $n$ can be readily fixed at the boundary along the time direction as described in [4], doing so at the geodesic’s other end in the bulk is not straightforward because of the non-covariance of bulk TMG, requiring a particular prescription. The normal vectors at the endpoints of the relevant geodesics is determined as follows. First, we choose an orthonormal triad $(v^{\mu},n^{\mu},\tilde{n}^{\mu})$ at the initial endpoint of interval where $v$ represents the tangent vector, $n$ is aligned with the boundary future time direction and $\tilde{n}$ is determined by orthogonality and the right hand convention. For the other endpoint of spacelike geodesic I as depicted in fig. 2, the vector $\tilde{n}$ is maintained as perpendicular to the plane (along $x$ direction) and $n$ is subsequently determined based on the tangent vector $v$ and $\tilde{n}$. Secondly, at the final endpoint of the interval, an orthonormal triad is established with $n$ remaining along the future time direction and $\tilde{n}$ derived from $n$ and $v$ which is now oriented towards the negative x direction. This specific $\tilde{n}$ is then utilized at the other endpoint of spacelike geodesic II to determine the normal vector $n$ using $v$ and $\tilde{n}$. Finally, these determined normal vectors at the endpoints of the spacelike geodesics in the bulk are smoothly connected across null infinity to find the normal vectors $n$ at the endpoints of the timelike geodesic.
For the spacelike geodesic I as shown in fig. 2 which starts from $t=-T/2$ at the boundary ($u\to 0$) and goes to infinity $u\to\infty$, it can be parameterised by $\beta$ as $t=-T/2\cosh\beta$ and $u=T/2\sinh\beta$ such that $\beta=0$ corresponds to the boundary and $\beta\to\infty$ for the other end point of the geodesic. The normal vector $n$ is timelike for the case of spacelike geodesic as required by condition (4.5). Since the other end point is at infinity and normal vector becomes null at $\beta\to\infty$, we regularize it by taking some finite but large value of $\beta$ so that it remains timelike near infinity and then take $\beta\to\infty$ at the end of computations. The boundary values of the (normalised) normal vector $n$ can be written as
$n^{i}_{\mu}=\frac{\ell}{u_{\epsilon}}(-1,0,0),\quad n^{f}_{\mu}=\frac{\ell}{u_{\infty}}(-\cosh\beta,0,\sinh\beta),$ (4.27)
where the initial value of normal vector at boundary is same as considered in [4] and determined by the boundary CFT data. The final value of $n$ is perpendicular to geodesic at null infinity (as the geodesic approaches the line $t=u$ at $u\to\infty$).
Similarly for curve II, we choose the boundary conditions on $n$ as follows
$n^{i}_{\mu}=\frac{\ell}{u_{\infty}}(-\cosh\beta,0,-\sinh\beta),\quad n_{\mu}^{f}=\frac{\ell}{u_{\epsilon}}(-1,0,0),$ (4.28)
where the curve II is parametrized by $\beta$ as $t=T/2\cosh\beta$ and $u=T/2\sinh\beta$ such that $\beta=0$ corresponds to the boundary and $\beta\to\infty$ denotes the other end point of geodesic II in the bulk. Note that these vectors are normalised with respect to the bulk metric.
Upon using (4.26), (4.27) and (4.28) in (4.11), we find that the total anomalous contribution from spacelike geodesics I and II vanishes i.e
$S_{\rm anom}^{\rm spacelike}=0.$ (4.29)
This is very similar to the case for the entanglement entropy for an unboosted interval (i.e a pure spacelike interval) [4] where anomalous contribution also vanishes. Therefore, the contribution from the spacelike geodesics to the on-shell action (4.6) is given by
$S_{A}^{\rm spacelike}=\frac{\ell}{2G_{N}^{(3)}}\log\frac{T}{\epsilon}.$ (4.30)
Timelike geodesic contribution
For the timelike geodesic (green color curve in fig. 3) as described by (4.24b), the length of this geodesic can be obtained as
$L^{\rm timelike}_{A}=2i\ell R\int_{R}^{\infty}\frac{du}{u}\frac{1}{\sqrt{u^{2}-R^{2}}}=i\ell\pi.$ (4.31)
Figure 3: Schematic for the boundary value of normal vectors $n_{1}$ and $n_{2}$ which are spacelike for timelike geodesics.
Now to find the anomalous contribution, the tangent vector and parallel transported vector to this timelike geodesic satisfying the condition (4.20) can be written as
$v^{\mu}=\frac{u}{\ell R}(u,0,t),\quad q^{\mu}=-\frac{u}{\ell R}(t,0,u),\quad\tilde{q}^{\mu}=(0,\frac{u}{\ell},0).$ (4.32)
The timelike geodesic (4.24b) can be parameterized by $\gamma$ as $u=R\cosh\gamma,t=R\sinh\gamma$ where $\gamma$ range from $\infty$- ∞ to +$\infty$+ ∞. Given that the timelike geodesic matches up with the spacelike geodesics I, II asymptotically at the null infinities, we propose to specify the initial and final normal vectors $n_{1}$, $n_{2}$ by matching them up with the normal vectors $n_{i}$ and $n_{f}$ at the respective null infinities (point 1 and 2 in fig. 3). As such we have
$n^{1}_{\mu}=\frac{\ell}{u_{-\infty}}(\sinh\gamma,0,\cosh\gamma),$ (4.33)
where $\gamma$ goes to $\infty$- ∞ near the initial point 1. As the endpoints extend to infinity, we introduce a regularization for computational purposes by choosing $\gamma$ a large value similar to the previous subsection. This ensures the boundary value of normal vectors remain spacelike near infinity i.e at the intersection of spacelike and timelike geodesics. Similarly we have
$n^{2}_{\mu}=\frac{\ell}{u_{\infty}}(-\sinh\gamma,0,-\cosh\gamma),$ (4.34)
with $\gamma\to\infty$ for the normal vector near the point 2 at the null infinity. Now using (4.32), (4.33) and (4.34) in (4.20), we obtain the anomalous contribution from timelike geodesic as
$S_{\rm anom}^{\rm timelike}$ $=\frac{1}{4G_{N}^{(3)}\mu}\log\left[\frac{q(s_{2})\cdot n_{2}-i\tilde{q}(s_{2})\cdot n_{2}}{q(s_{1})\cdot n_{1}-i\tilde{q}(s_{1})\cdot n_{1}}\right]$ (4.35)
$=\frac{1}{4G_{N}^{(3)}\mu}i\pi.$
Hence the total contribution to the on-shell action (4.15) for timelike geodesic can be obtained using (4.31) and (4.35) as
$S_{A}^{\rm timelike}=\frac{\ell}{4G_{N}^{(3)}}i\pi+\frac{1}{4G_{N}^{(3)}\mu}i\pi.$ (4.36)
Now upon using (4.30) and (4.36) in (4.22), we obtain the holographic TEE in the presence of gravitational anomaly after using (4.3) as follows
$S_{A}^{T}=\frac{c_{L}+c_{R}}{6}\log\frac{T}{\epsilon}+\frac{c_{R}}{6}i\pi.$ (4.37)
Interestingly, we observe that the above result matches exactly with the dual field theory result (3.5).
4.4 Holographic TEE at finite temperature and angular potential
We now compute the holographic timelike entanglement entropy for a pure timelike interval in a dual CFT2 at finite temperature and angular potential. The bulk dual for this case is described by the rotating BTZ black hole (black brane) with the metric
$ds^{2}=-\frac{(r^{2}-r_{+}^{2})(r^{2}-r_{-}^{2})}{r^{2}\ell^{2}}d\tau^{2}+\frac{r^{2}\ell^{2}}{(r^{2}-r_{+}^{2})(r^{2}-r_{-}^{2})}dr^{2}+r^{2}(d\phi+\frac{r_{+}r_{-}}{r^{2}\ell}d\tau)^{2}.$ (4.38)
where the coordinate $\phi$ is non-compact, and $r_{\pm}$ are the radii of the outer and inner horizons respectively. The left and right moving temperatures corresponding to the left and the right moving modes in a dual CFT2 can be written in terms of horizon radius as
$T_{L(R)}=\frac{1}{\beta_{L(R)}}=\frac{r_{+}\mp r_{-}}{2\pi\ell}.$ (4.39)
To obtain the geodesic length and anomalous contribution in this geometry, we can use the fact that the rotating BTZ black hole is locally equivalent to pure AdS3 so that one can map the metric (4.38) to the Poincaré patch of the AdS3 spacetime as follows
$x\pm t=\sqrt{\frac{r^{2}-r_{+}^{2}}{r^{2}-r_{-}^{2}}}e^{2\pi(\phi\pm\tau)/\beta_{R,L}},\,\,\qquad u=\sqrt{\frac{r_{+}^{2}-r_{-}^{2}}{r^{2}-r_{-}^{2}}}e^{(\phi r_{+}+\tau r_{-})/\ell}.$ (4.40)
Then the metric of AdS3 in the Poincaré coordinates is given by
$ds^{2}=\frac{\ell^{2}}{u^{2}}(-dt^{2}+dx^{2}+du^{2}).$ (4.41)
Now consider a pure timelike interval $A$ of length $T$ having end points $(\tau_{1},\phi_{1})=(-\frac{T}{2},0)$ and $(\tau_{2},\phi_{2})=(\frac{T}{2},0)$ in a dual CFT2 at finite temperature and angular potential. The corresponding end points of the interval $A$ in the Poincaré coordinates i.e $(t_{1},x_{1})$ and $(t_{2},x_{2})$ can be obtained using the map (4.40). The equations for the geodesic to compute the holographic TEE corresponding to this timelike interval can be written as
$\lambda^{2}-u^{2}$ $=\frac{R_{P}^{2}}{4},$ (4.42)
$u^{2}-\lambda^{2}$ $=R_{0}^{2}$
where $\lambda$ parametrizes the geodesics, mirroring the similar role $t$ played in (4.24). The first equation represents the spacelike geodesics joining end points of interval and null infinities while second equation is the timelike geodesic which joins these two null infinities. Here $R_{0}$ is some constant and $R_{P}$ is the length of timelike interval $A$ in the Poincaré coordinates as
$R_{P}=\sqrt{(t_{2}-t_{1})^{2}-(x_{2}-x_{1})^{2}}.$ (4.43)
Now, the total length of these geodesics in Poincaré coordinates can be obtained as
$L=\log\frac{R_{P}^{2}}{\epsilon_{1}\epsilon_{2}}+i\pi,$ (4.44)
where the relation between the cut-off $\epsilon_{i}$ ($i=1,2$) and $r_{\infty}$ for the bulk in the Poincaré coordinates and rotating BTZ respectively, can be written using (4.40) as
$\varepsilon_{i}=\sqrt{\frac{r_{+}^{2}-r_{-}^{2}}{r_{\infty}^{2}}}e^{\tau_{i}r_{-}/\ell}.$ (4.45)
Upon utilizing the coordinate transformation (4.40) and (4.45) in (4.44), we obtain the length contribution to holographic TEE in terms of the original rotating BTZ coordinates as
$S_{\rm EH}=\frac{1}{4G_{N}^{(3)}}\log\left(\frac{\beta_{L}\beta_{R}}{\pi^{2}\epsilon^{2}}\sinh\frac{\pi T}{\beta_{L}}\sinh\frac{\pi T}{\beta_{R}}\right)+\frac{1}{4G_{N}^{(3)}}i\pi,$ (4.46)
where $\epsilon$ is the UV cut-off for the dual CFT2 and related to the bulk infra red cut-off $r_{\infty}$ as $\epsilon={\ell}/{r_{\infty}}$.
For the anomalous contribution, we can compute it separately for contribution from spacelike and timelike geodesics, however doing it is very complicated in rotating BTZ background. To compute it more simply, we use the fact that geodesic comprising of spacelike and timelike geodesic forms a single continuous curve homologous to the boundary interval and the integral $\int_{C}\tilde{n}\cdot\nabla n$ is insensitive to variation of $n$ along the curve and, only depends on the initial and final value of normal vector $n$. Since the initial and final points of geodesics lie on the spacelike geodesic, the boundary conditions on the normal vector $n$ can be written in terms of time coordinate used to describe the black hole similar to [4]
$n_{i}=n_{f}=\left(\partial_{\tau}\right)_{\rm CFT}.$ (4.47)
The parallel transported vectors along the spacelike geodesics in the Poincaré coordinates can be obtained by boosting (4.26) as [^4] Note that the Lorentz boosts preserve the causal structure of the metric ensuring that timelike intervals (geodesics) remain timelike and spacelike intervals (geodesics) remain spacelike.
$q^{\mu}=-\frac{2u}{\ell R_{P}}\left(\frac{\lambda t_{21}}{R_{P}},\frac{\lambda x_{21}}{R_{P}},u\right),\,\,\quad\tilde{q}^{\mu}=\frac{u}{\ell}\left(\frac{x_{21}}{R_{P}},\frac{t_{21}}{R_{P}},0\right).$ (4.48)
Here $\lambda$ plays the role of $t$ as discussed earlier and takes value at initial point $\lambda_{i}=-\frac{R_{P}}{2}$ and $\lambda_{f}=\frac{R_{P}}{2}$ at the final point. Now using the coordinate transformation (4.40) to transform $(q,\tilde{q})$ into the BTZ coordinates followed by (4.47) and (4.11), we obtain the anomaly contribution to holographic TEE as follows
$S_{\rm anom}=\frac{1}{4G_{N}^{(3)}\mu}\log\left(\frac{\sinh\frac{\pi T}{\beta_{R}}\beta_{R}}{\sinh\frac{\pi T}{\beta_{L}}\beta_{L}}\right)+\frac{1}{4G_{N}^{(3)}\mu}i\pi.$ (4.49)
We observe that it is complex and the imaginary part is similar to the zero temperature case. The holographic TEE at finite temperature with a conserved angular momentum can now be written using (4.46), (4.49) and (4.3) as follows
$S_{A}^{T}=\frac{c_{L}+c_{R}}{12}\log\left(\frac{\beta_{L}\beta_{R}}{\pi^{2}\epsilon^{2}}\sinh\frac{\pi T}{\beta_{L}}\sinh\frac{\pi T}{\beta_{R}}\right)-\frac{(c_{L}-c_{R})}{12}\log\left(\frac{\beta_{R}\sinh\frac{\pi T}{\beta_{R}}}{\beta_{L}\sinh\frac{\pi T}{\beta_{L}}}\right)+\frac{c_{R}}{6}i\pi.$ (4.50)
We again see that the above result matches exactly with corresponding dual field theory result described in (3.7).
4.5 Holographic TEE at zero temperature and finite angular potential
We now turn our attention to compute the holographic TEE in CFT2 with a conserved charge which is dual to the extremal rotating BTZ black hole. The metric for the extremal BTZ black hole can be obtained by taking the extremal limit $(r_{+}=r_{-}=r_{0})$ in (4.38) as follows
$ds^{2}=-\frac{(r^{2}-r_{0}^{2})^{2}}{r^{2}\ell^{2}}d\tau^{2}+\frac{r^{2}\ell^{2}}{(r^{2}-r_{0}^{2})^{2}}dr^{2}+r^{2}(d\phi+\frac{r_{0}^{2}}{r^{2}\ell}d\tau)^{2}.$ (4.51)
In the extremal limit, the Hawking temperature $T_{H}$ and left moving temperature $T_{L}$ vanishes but there remains an effective temperature called the Frolov-Thorne temperature that represents the degeneracy of the ground state of an extremal BTZ black hole and given by [67]
$T_{FT}=\frac{r_{0}}{\pi\ell}=\frac{1}{\beta_{R}}.$ (4.52)
The metric in (4.51) can be mapped to the pure AdS3 spacetime in the Poincaré coordinates (4.41) by the following coordinate transformations [74]
$x+t$ $=\frac{\ell}{2r_{0}}e^{\frac{2r_{0}}{\ell}(\phi+\tau)},\quad\quad x-t=\phi-\tau-\frac{\ell r_{0}}{r^{2}-r_{0}^{2}}$ (4.53)
$u$ $=\frac{\ell}{\sqrt{r^{2}-r_{0}^{2}}}e^{r_{0}(\phi+\tau)/\ell},$
Consider a pure timelike interval $A$ of length $T$ having end points $(\tau_{1},\phi_{1})=(-\frac{T}{2},0)$ and $(\tau_{2},\phi_{2})=(\frac{T}{2},0)$ in a dual CFT2 at zero temperature and finite angular potential. The computation of the length of the spacelike and timelike geodesic is similar to the previous subsection involving the non-extremal black hole. So, the contribution from length of geodesics to the holographic TEE for this case can be obtained using (4.53) in (4.44) as follows
$S_{\rm EH}=\frac{1}{4G_{N}^{(3)}}\log\left(\frac{\ell T}{r_{0}\epsilon^{2}}\sinh\frac{r_{0}T}{\ell}\right)+\frac{1}{4G_{N}^{(3)}}i\pi,$ (4.54)
where $\epsilon$ is the UV cut-off. Similarly we can obtain the anomalous contribution by transforming the parallel transported vectors (4.48) using (4.53) and subsequently using (4.47) in (4.11) which is given by
$S_{\rm anom}=\frac{1}{4G_{N}^{(3)}\mu}\log\left(\frac{\ell\sinh\frac{r_{0}T}{\ell}}{r_{0}T}\right)+\frac{1}{4G_{N}^{(3)}\mu}i\pi.$ (4.55)
The holographic TEE at zero temperature with a conserved angular momentum may now be obtained upon using (4.54), (4.55) and (4.3) as follows
$S_{A}^{T}$ $=\frac{c_{L}+c_{R}}{12}\log\left(\frac{T\,\ell}{r_{0}\epsilon^{2}}\sinh\frac{r_{0}T}{\ell}\right)-\frac{(c_{L}-c_{R})}{12}\log\left(\frac{\ell\sinh\frac{r_{0}T}{\ell}}{r_{0}T}\right)+\frac{c_{R}}{6}i\pi$ (4.56)
$=\frac{c_{L}}{6}\log\frac{T}{\epsilon}+\frac{c_{R}}{6}\log\left(\frac{\ell}{r_{0}\epsilon}\sinh\frac{r_{0}T}{\ell}\right)+\frac{c_{R}}{6}i\pi.$
We observe that the imaginary part corresponding to left moving mode is absent similar to earlier results and terms containing $c_{R}$ resembles the TEE at an effective Frolov-Thorne temperature [67]. It also matches with the dual field theory result given in (3.8).
5 Summary and discussion
To summarize, we have studied the effect of gravitational anomaly on the timelike entanglement entropy (TEE) in the context of AdS3/CFT2 duality. Specifically we employed analytic continuation and computed the TEE for a pure timelike interval in a CFT2 with gravitational anomaly in the vacuum state of CFT2, for thermal CFT2’s with an angular potential and also for CFT2’s at finite angular potential with zero temperature. As a result, the TEE receives an additional purely imaginary contribution due to the gravitational anomaly. Interestingly for a pure timelike interval in all cases, the imaginary part depends on only one central charge in contrast to its real counterpart which depends on both the left and right central charges. This chiral sensitivity of the TEE suggests that it may be used as a probe for gravitational anomalies. We see that in the absence of the gravitational anomaly, our results reduce to the known results in the CFT2.
Subsequently we computed the timelike entanglement entropy for a pure timelike interval in CFT2 with gravitational anomalies dual to bulk topologically massive gravity (TMG) in asymptotically AdS3 geometries. The bulk three dimensional action for the TMG-AdS3 geometries involve a gravitational Chern-Simons term in addition to the usual Einstein-Hilbert term. In the locally AdS3 geometry we considered, the worldlines of massive spining particles are geodesics. Since the holographic TEE consists of spacelike and timelike geodesics, it involves the on-shell actions of massive spinning particles moving on these extremal worldlines (geodesics) in the dual bulk geometry. We derived from the Chern-Simons action the anomalous contribution specifically arising from timelike geodesic which represents the rotation of normal frames along it. Therefore, the holographic TEE involves the sum of length of the geodesics and the twist (rotation) of the normal frame along these geodesics. We further proposed a prescription for the boundary condition on normal vectors to compute the Chern-Simons contribution to the holographic TEE. Utilizing this prescription, we determined the holographic TEE in a CFT2 dual to pure AdS3 and also for the cases of rotating extremal and non-extremal BTZ black holes. Interestingly we observed that the holographic results in all these cases agrees precisely with the dual field theory results.
There are several interesting directions to follow from our work. In this paper, we obtained the holographic TEE using spacelike and timelike geodesics. Exploring the same using the RT surface in a complexified geometry as proposed in [31] can offer new insights about the anomaly contribution. Furthermore, exploring TEE in the Chern-Simons formulation of $(2+1)$-dimensional topologically massive gravity theories [75] and geometries that deviate from locally AdS3 such as warped AdS [76] would be an interesting areas of investigation. Computing TEE in higher dimensional theories with gravitational anomalies will also be interesting. Finally it would be also interesting to investigate TEE in the context of boundary Lifshitz field theory [77] along the lines of AdS/BCFT [39]. These investigations can provide further new insights into black hole interiors and bulk reconstruction. We leave these interesting issue for future investigations.
Acknowledgments
We thank Dimitrios Giataganas for helpful discussions. H.P would like to thank Gaurav Katoch, Debajyoti Sarkar and Bhim Sen for collaboration during the initial stages of this work. C.S.C acknowledge support of this work by NCTS and the grant 113-2112-M-007-039-MY3 of the National Science and Technology Council of Taiwan. H.P acknowledges the support of this work by NCTS.